OpenStudy (anonymous):

using compatible numbers what is 813 divided by 30? how would you show your work for that?

OpenStudy (anonymous):

I think the best way would be to split 813 up into numbers easy to divide by 30: \[813=600+210+3\] \[600\div30=20\] \[210\div30=7\] \[3\div30=\frac{1}{10} \] So \(813 \div 30= 20+7+\frac{1}{10}\) Does this make sense to you?

OpenStudy (anonymous):

That does make sense but inclass we had to show our work like with the paper and pencil dividing way. how owuld you show your work like that? and the number has to be close to 813 that goes into 30 evenly

OpenStudy (anonymous):

Then just take an estimate by dividing 810 by 30, which is 27.

OpenStudy (anonymous):

To give a full answer, \(813\approx 810=600+210\) \(810\div 30 = (600\div 30)+(210\div 30)\) so \(810 \div 30 = 20 + 7 = 27\) Finally, \(813 \div 30 \approx 27\)

OpenStudy (anonymous):

okay I have one more question how do yous how your work with the paper and pencil way using compatible numbers for 2000 divided by 59